OpenStudy (anonymous):

X+10 /(X-2) (X+4) =A/X-2+B/X+4

OpenStudy (anonymous):

\[\frac{x+10}{(x-2)(x+4)}=\frac{A}{x-2}+\frac{B}{x-4}\] is this what you want?

OpenStudy (anonymous):

YEA LOL THANKS

OpenStudy (anonymous):

oops that is backwards. sorry

OpenStudy (anonymous):

let me do that again i had it backwards

OpenStudy (anonymous):

THE 4 ARE + 4 NOT _

OpenStudy (anonymous):

you need the method or just the answer?

OpenStudy (anonymous):

ok let me try one more time

OpenStudy (anonymous):

METHOD AND ANSWER PLZ

OpenStudy (anonymous):

ok first the answer \[A=2,B=1\]

OpenStudy (anonymous):

now method. if you were going to add \[\frac{A}{x-2}+\frac{B}{x+4}\] your numerator would be \[A(x+4)+B(x-2)\] and you know your numerator is \[x+10\] so set \[(x+4)A+(x-2)B=x+10\]

OpenStudy (anonymous):

this has to be an identity for all x so you get to pick any x to solve for A and B. put \[x=2\] gete \[(2+4)A+(2-2)B=2+10\] \[6A=12\] \[A=2\]

OpenStudy (anonymous):

YOUR SMART THAMKS

OpenStudy (anonymous):

now guess what you pick for x? pick \[x=-4\] get \[(-4+4)A+(-4-2)B=-4+10\] \[-6B=6\]\[B=-1\]